OpenStudy (anonymous):

need some more help lol find the slope for the following line; Y=3 OVER 4X-7 A; -7 B: 7 C: 3 0ver 4 D. -3 over 4 @Rhystic

OpenStudy (rhystic):

so you have y = 3/(4x-7) is that correct?

OpenStudy (anonymous):

its y-3/4x-7 but its 3/4 as in a fraction not like 3/4ths some people get confused @Rhystic

OpenStudy (rhystic):

oh that makes it a lot simpler. the equation for a straight line is y = mx + b where m is the slope and b is the y intercept. in your case you have y = (3/4)x - 7 which means m =.... ?

OpenStudy (anonymous):

i think its -4/3

OpenStudy (rhystic):

is the x in the denominator or the numerator?

OpenStudy (anonymous):

right sorry D is suppose to be -4 over 3

OpenStudy (anonymous):

i had to look over my options and i accidentaly flipped the 2 i think its -4/3

OpenStudy (rhystic):

where is x located? is it (3x/4) or 3/(4x) ? or is it 3/(4x-7) ?

OpenStudy (anonymous):

y=3/4x-7

OpenStudy (anonymous):

@Rhystic

OpenStudy (rhystic):

then yea youre right

OpenStudy (anonymous):

awesome can you help with a few more @Rhystic ?